# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 7.2 Convergence in 3D using physically relevant parameters
Next we investigate the behaviour of the proposed numerical methods in a 3D setting and taking model parameters more closely related to applications in tissue poroelasticity. We still use manufactured solutions to assess the accuracy of the formulation, but take an exact displacement that satisfies $\mathop{\rm div}\nolimits\boldsymbol{u}\to 0$ as $\lambda\to\infty$. The domain is the 3D box $\Omega=(0,L)\times(0,L)\times(0,2L)$ with $L=0.01$ m, and mixed boundary conditions were taken analogously as before, separating the domain boundary between $\Gamma_{\mathrm{N}}$ defined as the planes $x=0$, $y=0$ and $z=0$, and $\Gamma_{\mathrm{D}}$ as the remainder of the boundary. The manufactured displacement and pressure head are
$\boldsymbol{u}(x,y,z)=\frac{L}{4}\begin{pmatrix}\sin(x/L)\cos(y/L)\sin(z/(2L))+x^{2}/\lambda\\ -2\cos(x/L)\sin(y/L)\cos(z/(2L))+y^{2}/\lambda\\ 2\cos(x/L)\cos(y/L)\sin(z/(2L))-2z^{2}/\lambda\end{pmatrix},\quad p(x,y,z)=\sin(x/L)\cos(y/L)\sin(z/(2L)).$
DoFs $h$ $e(\boldsymbol{\sigma})$ rate $e(\boldsymbol{u})$ rate $e(\boldsymbol{\rho})$ rate $e({\boldsymbol{\varphi}})$ rate $e(p)$ rate
@@
First we set again the model parameters to mild values $\lambda=\mu=c_{0}=k_{0}=\alpha=\mu_{f}=1$, $k_{1}=k_{2}=0.1$, and we compare them against the following values (from, e.g., [9, 43])
$k_{0}=2.28\times 10^{-11}\,\text{m}^{3},\quad k_{1}=5\times 10^{-12}\,\text{m}^{3},\quad\lambda=1.44\times 10^{6}\,\text{Pa},\quad\mu=9.18\times 10^{3}\,\text{Pa},\quad\mu_{f}=7.5\times 10^{-4}\,\text{Pa}\cdot\text{s},$
and $c_{0}=0$, $\alpha=0.99$. Table 7.3 reports on the convergence of the method. While the magnitude of the stress errors is much higher for the second parameter regime, the discharge flux error magnitude is smaller than in the first case and the displacement, rotation, and fluid pressure errors remain roughly of the same magnitude. In any case, the table confirms that the optimal slope of the error decay is not affected by a vanishing storativity nor large Lamé constants. The numerical solutions are displayed in Figure 7.2.
@@ 7.3 Convergence in the case of adaptive mesh refinement
$\boldsymbol{u}(r,\theta)=\frac{r^{\chi}}{2\mu}\begin{pmatrix}-(\chi+1)\cos([\chi+1]\theta)+(M_{2}-\chi-1)M_{1}\cos([\chi-1]\theta)\\ (\chi+1)\sin([\chi+1]\theta)+(M_{2}+\chi-1)M_{1}\sin([\chi-1]\theta)\end{pmatrix},\qquad p(r,\theta)=r^{1/3}\sin\biggl{(}\frac{1}{3}(\frac{\pi}{2}+\theta)\biggr{)},$
with polar coordinates $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$, $\theta=\arctan(x_{2},x_{1})$, $\chi\approx 0.54448373$, $M_{1}=-\cos([\chi+1]\omega)/\cos([\chi-1]\omega)$, and $M_{2}=2(\lambda+2\mu)/(\mu+\lambda)$. The boundary conditions (taking as $\Gamma_{\mathrm{N}}$ the segments at $x=\pm 1$ and $y=\pm 1$ and $\Gamma_{\mathrm{D}}$ the remainder of the boundary) and forcing data are constructed from these solutions and the model parameters are $\lambda=10^{3}$, $\mu=10$, $k_{0}=\frac{1}{2}$, $\mu_{f}=c_{0}=k_{1}=0.1$, $\alpha=\frac{1}{4}$, where for this test we consider a Kozeny–Carman permeability form. As in [20], a sub-optimal rate of convergence is expected for the mixed elasticity sub-problem in its energy norm. Note that since the exact pressure is in $\mathrm{H}^{4/3-\epsilon}(\Omega)$ for any $\epsilon>0$ (cf. [37, Chapter 5]), it is still regular enough to have optimal convergence. However its gradient (and therefore also the exact discharge flux ${\boldsymbol{\varphi}}$) has a singularity located at the reentrant corner and therefore we expect an order of convergence of approximately $O(h^{1/3})$.
DoFs $h$ $e(\boldsymbol{\sigma})$ rate $e(\boldsymbol{u})$ rate $e(\boldsymbol{\rho})$ rate $e({\boldsymbol{\varphi}})$ rate $e(p)$ rate $\texttt{eff}(\Xi)$
@@
The numerical results of this test are reported in Table 7.4. We observe the expected sub-optimal convergence under an uniform mesh refinement while the optimal convergence in all variables is attained as the mesh is locally refined (the first three rows are very similar as most of the elements are refined in the first three steps. This can be controlled by the bulk density, here taken as $\zeta=9.5\cdot 10^{-5}$). We also note that the individual errors are approximately of the same magnitude in the last row of each section of the table, but for the adaptive case this is achieved using approximately 5.5% of the number of degrees of freedom needed in the uniformly refined case. The last column of the table again confirms the reliability and efficiency of the a posteriori error estimator. Note that for this case we compute the divergence part of the error norm in the stress and fluxes as projections of the momentum and mass residuals onto the displacement and pressure discrete spaces, respectively. We plot in Figure 7.3 the approximate displacement and pressure as well as sample triangulations obtained after a few adaptive refinement steps that confirm the expected agglomeration of vertices near the reentrant corner.
@@ 7.4 Adaptive computation of cross-sectional flow and deformation in a soft tissue specimen
Finally, we apply the proposed methods to simulate the localisation of stress, deformation, and flow patterns in a multi-layer cross-section of cervical spinal cord. We follow the setup in [42, 44]. The geometry and unstructured mesh have been generated using GMSH [35] from the images in [44]. The heterogeneous porous material consists of white and grey matter surrounded by the pia mater (a thin layer, also considered poroelastic. See Figure 7.4, top two left panels). All components are assumed fully saturated with cerebospinal fluid. The transversal cross-section has 1.3 cm in maximal diameter and the indentation region is a curved subset of the anterior part of the pia mater (a sub-boundary of length 0.4 cm). Boundary conditions are of mixed load-traction type, but slightly different than the ones analysed in the previous section. We conduct an indentation test applying a traction $\boldsymbol{\sigma}\mbox{\boldmath{$n$}}=(0,-P)^{-\tt t}$, with $P$ a constant solid pressure of 950 dyne/cm2. The posterior part of the pia mater acts as a rigid posterior support where we prescribe zero displacement. The remainder of the boundary of the pia mater is stress-free. For the fluid phase we impose a constant inflow pressure of cerebrospinal fluid of 1.1 dyne/cm2 and zero outflow pressure at the stress-free sub-boundary, as well as zero normal discharge flux at the posterior support. For the three different layers of the domain we use the following values for Young modulus, Poisson ratio, and lower bound for permeability (some values from [10, 42, 44]) $E^{\mathrm{pia}}=23^{\prime}000\,\text{dyne/cm}^{2},\ \nu^{\mathrm{pia}}=0.3$, $E^{\mathrm{white}}=8^{\prime}400\,\text{dyne/cm}^{2},\ \nu^{\mathrm{white}}=0.479,\ E^{\mathrm{grey}}=16^{\prime}000\,\text{dyne/cm}^{2}$, $\nu^{\mathrm{grey}}=0.49$, $k_{0}^{\mathrm{grey}}=1.4\cdot 10^{-9}\,\text{dyne/cm}^{2}$, $k_{0}^{\mathrm{white}}=1.4\cdot 10^{-6}\,\text{dyne/cm}^{2}$, $k_{0}^{\mathrm{pia}}=3.9\cdot 10^{-10}\,\text{dyne/cm}^{2}$. Further, we take $\boldsymbol{f}=\boldsymbol{0}$, $g=0$, $\mu_{f}=70\,\text{dyne/cm}^{2}\cdot$ s (for cerebrospinal fluid at 37), $k_{1}=\frac{1}{2}k_{0}$, $\alpha=\frac{1}{4}$, and $c_{0}=10^{-3}$.

### v2

@@ 7.2 Convergence in 3D using physically relevant parameters
Next we investigate the behaviour of the proposed numerical methods in a 3D setting and taking model parameters more closely related to applications in tissue poroelasticity. We still use manufactured solutions to assess the accuracy of the formulation, but take an exact displacement that satisfies $\mathop{\rm div}\nolimits{\bm{u}}\to 0$ as $\lambda\to\infty$. The domain is the 3D box $\Omega=(0,L)\times(0,L)\times(0,2L)$ with $L=0.01$ m, and mixed boundary conditions were taken analogously as before, separating the domain boundary between $\Gamma_{\mathrm{N}}$ defined by the planes $x=0$, $y=0$ and $z=0$, and $\Gamma_{\mathrm{D}}$ as the remainder of the boundary. The manufactured displacement and pressure head are
${\bm{u}}=\frac{L}{4}\begin{pmatrix}\sin(x/L)\cos(y/L)\sin(z/(2L))+x^{2}/\lambda\\ -2\cos(x/L)\sin(y/L)\cos(z/(2L))+y^{2}/\lambda\\ 2\cos(x/L)\cos(y/L)\sin(z/(2L))-2z^{2}/\lambda\end{pmatrix},\quad p=\sin(x/L)\cos(y/L)\sin(z/(2L)).$
DoFs $h$ $e(\bm{\sigma})$ rate $e({\bm{u}})$ rate $e(\bm{\rho})$ rate $e({\bm{\varphi}})$ rate $e(p)$ rate
@@
First we set again the model parameters to mild values $\lambda=\mu=c_{0}=k_{0}=\alpha=\mu_{f}=1$, $k_{1}=k_{2}=0.1$, we use the exponential permeability constitutive law, and we compare them against the following values (from, e.g., [9, 44])
$k_{0}=2.28\times 10^{-11}\,\text{m}^{3},\quad k_{1}=5\times 10^{-12}\,\text{m}^{3},\quad\lambda=1.44\times 10^{6}\,\text{Pa},$
$\mu=9.18\times 10^{3}\,\text{Pa},\quad\mu_{f}=7.5\times 10^{-4}\,\text{Pa}\cdot\text{s},\quad c_{0}=0,\quad\alpha=0.99.$
Table 7.3 reports on the convergence of the method. While the magnitude of the stress errors is much higher for the second parameter regime, the discharge flux error magnitude is smaller than in the first case and the displacement, rotation, and fluid pressure errors remain roughly of the same magnitude. In any case, the table confirms that the optimal slope of the error decay is not affected by a vanishing storativity nor large Lamé constants. The components of the numerical solution are displayed in Figure 7.2.
@@ 7.3 Convergence in the case of adaptive mesh refinement
${\bm{u}}(r,\theta)=\frac{r^{\chi}}{2\mu}\begin{pmatrix}-(\chi+1)\cos([\chi+1]\theta)+(M_{2}-\chi-1)M_{1}\cos([\chi-1]\theta)\\ (\chi+1)\sin([\chi+1]\theta)+(M_{2}+\chi-1)M_{1}\sin([\chi-1]\theta)\end{pmatrix},$
$p(r,\theta)=r^{1/3}\sin\biggl{(}\frac{1}{3}(\frac{\pi}{2}+\theta)\biggr{)},$
with polar coordinates $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$, $\theta=\arctan(x_{2},x_{1})$, $\chi\approx 0.54448373$, $M_{1}=-\cos([\chi+1]\omega)/\cos([\chi-1]\omega)$, and $M_{2}=2(\lambda+2\mu)/(\mu+\lambda)$. The boundary conditions (taking as $\Gamma_{\mathrm{N}}$ the segments at $x=\pm 1$ and $y=\pm 1$ and $\Gamma_{\mathrm{D}}$ the remainder of the boundary) and forcing data are constructed from these solutions and the model parameters are $\lambda=10^{3}$, $\mu=10$, $k_{0}=\frac{1}{2}$, $\mu_{f}=c_{0}=k_{1}=0.1$, $\alpha=\frac{1}{4}$, where for this test we consider a Kozeny–Carman permeability form. As in [20], a sub-optimal rate of convergence is expected for the mixed elasticity sub-problem in its energy norm. Note that since the exact pressure is in $\mathrm{H}^{4/3-\epsilon}(\Omega)$ for any $\epsilon>0$ (cf. [38, Chapter 5]), it is still regular enough to have optimal convergence. However its gradient (and therefore also the exact discharge flux ${\bm{\varphi}}$) has a singularity located at the reentrant corner and therefore we expect an order of convergence of approximately $O(h^{1/3})$.
DoFs $h$ $e(\bm{\sigma})$ rate $e({\bm{u}})$ rate $e(\bm{\rho})$ rate $e({\bm{\varphi}})$ rate $e(p)$ rate $\texttt{eff}(\Xi)$
@@
The numerical results of this test are reported in Table 7.4. We observe the expected sub-optimal convergence under an uniform mesh refinement while the optimal convergence in all variables is attained as the mesh is locally refined (the first three rows are very similar as most of the elements are refined in the first three steps. This can be controlled by the bulk density, here taken as $\zeta=9.5\cdot 10^{-5}$). We also note that the individual errors are approximately of the same magnitude in the last row of each section of the table, but for the adaptive case this is achieved using approximately 5.5% of the number of degrees of freedom needed in the uniformly refined case. The last column of the table again confirms the reliability and efficiency of the a posteriori error estimator. Note that for this case we compute the divergence part of the error norm in the stress and fluxes as projections of the equilibrium and mass residuals onto the displacement and pressure discrete spaces, respectively. We plot in Figure 7.3 the approximate displacement and pressure as well as sample triangulations obtained after a few adaptive refinement steps that confirm the expected agglomeration of vertices near the reentrant corner.
@@ 7.4 Adaptive computation of cross-sectional flow and deformation in a soft tissue specimen
Finally, we apply the proposed methods to simulate the localisation of stress, deformation, and flow patterns in a multi-layer cross-section of cervical spinal cord. We follow the setup in [43, 45]. The geometry and unstructured mesh have been generated using GMSH [36] from the images in [45]. The heterogeneous porous material consists of white and grey matter surrounded by the pia mater (a thin layer, also considered poroelastic. See Figure 7.4, top two left panels). All components are assumed fully saturated with cerebrospinal fluid. The transversal cross-section has 1.3 cm in maximal diameter and the indentation region is a curved subset of the anterior part of the pia mater (a sub-boundary of length 0.4 cm). Boundary conditions are of mixed load-traction type, but slightly different than the ones analysed in the previous section. We conduct an indentation test applying a traction $\bm{\sigma}\mbox{\boldmath{$n$}}=(0,-P)^{-\tt t}$, with $P$ a constant solid pressure of 950 dyne/cm2. The posterior part of the pia mater acts as a rigid posterior support where we prescribe zero displacement. The remainder of the boundary of the pia mater is stress-free. For the fluid phase we impose a constant inflow pressure of cerebrospinal fluid of 1.1 dyne/cm2 and zero outflow pressure at the stress-free sub-boundary, as well as zero normal discharge flux at the posterior support. For the three different layers of the domain we use the following values for Young modulus, Poisson ratio, and lower bound for permeability (some values from [10, 43, 45]) $E^{\mathrm{pia}}=23^{\prime}000\,\text{dyne/cm}^{2},\ \nu^{\mathrm{pia}}=0.3$, $E^{\mathrm{white}}=8^{\prime}400\,\text{dyne/cm}^{2},\ \nu^{\mathrm{white}}=0.479,\ E^{\mathrm{grey}}=16^{\prime}000\,\text{dyne/cm}^{2}$, $\nu^{\mathrm{grey}}=0.49$, $k_{0}^{\mathrm{grey}}=1.4\cdot 10^{-9}\,\text{dyne/cm}^{2}$, $k_{0}^{\mathrm{white}}=1.4\cdot 10^{-6}\,\text{dyne/cm}^{2}$, $k_{0}^{\mathrm{pia}}=3.9\cdot 10^{-10}\,\text{dyne/cm}^{2}$. Further, we take $\bm{f}=\bm{0}$, $g=0$, $\mu_{f}=70\,\text{dyne/cm}^{2}\cdot$ s (for cerebrospinal fluid at 37), $k_{1}=\frac{1}{2}k_{0}$, $\alpha=\frac{1}{4}$, and $c_{0}=10^{-3}$.
@@
8 Concluding remarks
In this work, we developed a family of mixed finite element methods for a nonlinear poroelasticity model with stress-dependent permeability. By reformulating the constitutive equations, we enabled the use of a Hellinger–Reissner-type mixed formulation, ensuring robustness with respect to near-incompressibility and vanishing storativity limits. The problem structure was analysed using fixed-point theory and saddle-point formulations, leading to well-posedness results for both the continuous and discrete settings. The chosen finite element spaces (PEERSk elements for elasticity and Raviart–Thomas elements for fluid flow) provided exact equilibrium and mass conservation.
We derived a priori error estimates, demonstrating optimal convergence properties. Furthermore, a residual-based a posteriori error estimator was introduced and shown to be both reliable and efficient. This estimator guided adaptive mesh refinement strategies, which were validated through numerical experiments in both two and three dimensions.
The proposed nonlinear dependence of permeability on stress in poroelasticity is particularly relevant for soft tissues, where permeability changes due to deformation, and for subsurface reservoirs, where stress variations influence fluid transport. The current formulation can be adapted to different material systems by choosing appropriate functional forms for the permeability-stress relationship including also anisotropic permeability variations.
We are keen to follow other extensions, for instance exploring different abstract results that would impose a less restrictive assumption on smallness of data, and extend the formulation to accommodate a nonlinear stress-strain constitutive law, transient effects, as well as fully coupled multiphysics systems, such as thermo-poroelasticity. Note however that these generalisations (useful to broaden the applicability of the proposed methods to more complex real-world scenarios) will require a substantially different theoretical framework.
Funding
This work has been supported in part by the Australian Research Council through the Future Fellowship grant FT220100496 and Discovery Project grant DP22010316; by the National Research and Development Agency (ANID) of the Ministry of Science, Technology, Knowledge and Innovation of Chile through the postdoctoral program Becas Chile grant 74220026.